OpenStudy (anonymous):

classify the system of equations. 2x=5-y -3+y=-2x+3

OpenStudy (anonymous):

By "classifying" do you just mean solve?

OpenStudy (anonymous):

1st eq=> 2x+y=5 2nd eq=> 2x+y=6 so adding them up you hhave...=>4x+2y=11 subtraction give 0=0